C&A Printing takes 10 seconds to print a page in color, but after 500 pages, the printer must be cooled down for 15 minutes. No pages can be printed while the printer is cooling down. What is C&A's capacity in pages per hour?
 
C&A's demand rate is 3 pages per minute. What is the smallest batch size to ensure the process is demand-constrained?
Answer #1

Answer 7:

  • Processing time = A =10 secs
  • Cooling time = B = 15 mins = 900 secs
  • Total time required to pring 500 pages = (A*500)+B= 5900 secs = 5900/3600 = 1.64 hours
  • So hourly capacity = 500 / total time required to make 500 = 500/1.64
  • Hourly capacity = 305.0847 pages per hour

3page per minute :

  • Let the minimum no. of pages for 3 page per minute cnstraint be X.
  • Total time for producing X = (A*X) + B= AX+900
  • So average per unit time = total time for X / X = (10X+900)/X
  • But average per unit time required = 1/ 3 mins = 20 Secs (Considering 3 pages per min)
  • Equating both :
    • (10X+900)/X = 20
    • 10X+900 = 20X
    • 10X = 900
    • X= 90 (Answer)
  • So the smallest lot size to satisfy 3 pages per minute constraint = 90 pages per lot